what is the center (h,k) of 3x^2+12x+3y^2=0
Answer by nerdybill(7384)

what is the center (h,k) of 3x^2+12x+3y^2=0
.
put in standard form of circle:
(x-h)^2 + (y-k)^2 = r^2
.
3x^2+12x+3y^2=0
(3x^2+12x)+(3y^2)=0
3(x^2+4x) + 3(y^2)=0
3(x^2+4x+4) + 3(y^2)=12
3(x+2)^2 + 3(y^2) = 12
3(x-(-2))^2 + 3(y-0)^2 = 12
(x-(-2))^2 + (y-0)^2 = 4
center: (h,k) = (-2,0)